If 7 lb. of tea and 5 lb. of coffee cost $55.00, and 6 lb. of tea and 3 lb. of coffee cost $42.00, what is the price per pound of each?
Answer by stanbon(57347)

If 7 lb. of tea and 5 lb. of coffee cost $55.00, and 6 lb. of tea and 3 lb. of coffee cost $42.00, what is the price per pound of each?
---
Value Eq::: 7t + 5c = 55
Value Eq::: 6t + 3c = 42
------
Multiply thru 1st eq. by 3 and 2nd eq by 5:
21t + 15c = 3*55
30t + 15c = 5*42
-----------------------
Subtract 1st from 2nd and solve for "t":
9t = 5*42-3*55
3t = 5*14- 55
t = $5 (price of tea per pound)
---
Substitute into 6t+3c = 42 to solve for "c":
6*5 + 3c = 42
3c = 12
c = $4 (price of coffee per pound)
======================================
Cheers,
Stan H.